XxDoogxX:

Carter pushes a bag full of basketball jerseys across the gym floor. Then he pushes with a constant force of 21 newtons. If he pushes the bag 9 meters in 3 seconds, how much power does he use? (Hint: 1 watt = 1 .) A. 3 watts B. 7 watts C. 27 watts D. 63 watts E. 189 watts

SemiDefinite:

Careful. You don't *use* power. Power is the instantaneous rate at which energy +/or work is supplied/ applied. If i am doing a bench-press, I am doing work and using energy, but power is just a metric that measures all of this against time. Work is like Energy, and is measured in Joules (J): - W = F * s Power [units: W, Watts = J per sec] is about the delivery of energy (J), or work (J), per unit time (sec): - P = W/t (J/sec) Here, you have: 1) F = 21N, 2) s = 9m 3) t = 3 sec He does work W = 21 [N] * 9 [m] = 189[ J] That's the work done. He did it in t = 3 sec So the average Power over that t = 3 sec period of time is \(\dfrac{ 189}{3} = 63 W\)